# A Closer Look at Knowledge Distillation in Spiking Neural Network Training

### v1

@@ Introduction
Results on ImageNet-1K. The comparison results are presented in Tab. 2. The proposed CKDSNN continues to outperform previous SOTA methods using various teacher models, including ResNet-18, ResNet-34, and ResNet-50. Specifically, compared to the prior SOTA methods BKDSNN (xu2025bkdsnn), our CKDSNN improves the top-1 accuracy by 1.32%, 1.81%, and 1.52% using the three types of network architectures, respectively. These results indicate that CKDSNN is effective on the large-scale dataset.
Results on CIFAR10-DVS. The neuromorphic dataset comparison results are shown in Tab. 4. Our CKDSNN also outperforms existing methods on the neuromorphic dataset. For example, under the same architecture and time step settings, CKDSNN achieves an accuracy improvement of 1.05% compared to the most competitive EnOFSNN (guo2024enof).
@@
STBP (wu2019direct) NIPS’21 ResNet19 4 67.80
SEW-R (fang2021sew) NIPS’21 WideNet 16 74.40
KDSNN (Xu2023kdsnn) CVPR’23 ResNet20 10 78.31
BKDSNN (xu2025bkdsnn) ECCV’24 ResNet20 10 79.53
EnOFSNN (guo2024enof) NIPS’24 ResNet20 10 80.50
@@
Importance of CAM-SAM distillation in SAMD. The CAM-SAM distillation is one of the core components of SAMD. As show in Fig. 4 (b), when using ANN-based strategies (e.g., e2KD (parchami2024good) or CATKD (guo2023class)) to distill SNNs, the performance is significantly lower than our proposed CAM-SAM distillation. Specifically, CAM-CAM methods use class Grad-CAM strategies to generate activation maps, but the gradient estimation error leads to performance degradation. Additionally, CATKD is only applicable to CNN-base architectures. In contrast, SAM is designed on the characteristics of SNNs, enabling compatibility with various architectures.
@@
Figure 7: The t-SNE (van2008visualizing) visualization of feature distributions. (a): teacher ANN’s float feature representation. (b)-(d): spike features of student SNNs. Each color denotes an image class.
@@
Spiking Features. We extract spiking features from the final layer of our CKDSNN model and visualize them using t-SNE (van2008visualizing). Image samples from the CIFAR-100 dataset are used. As shown in Fig. 7, the spike features generated by CKDSNN demonstrate improved separability and discriminability compared to other KD methods and are close to the teacher’s feature distribution. This superiority is largely contributed to the proposed activation map KD, which guides the spiking representations to focus on salient regions, resulting in better feature discrimination.
@@
SEW-R (fang2021sew) 18.0 4.14 4.03 67.04 4
KDSNN (Xu2023kdsnn) 16.0 4.13 4.01 67.18 4
BKDSNN (xu2025bkdsnn) 15.0 4.02 3.98 71.24 4
@@
Energy Efficiency. We finally evaluate the energy efficiency of our method by calculating the Fire rate, SOPS, and Power consumption following prior methods (zhou2023spikformer; xu2025bkdsnn; fang2023spikingjelly). As shown in Tab. 6, compared with prior KD methods, our CKDSNN not only achieves state-of-the-art performance but offers higher energy efficiency. For example, our method with the same 4 time steps outperforms BKDSNN (xu2025bkdsnn) by 1.81% but using slightly less fire rate and Power. The energy efficiency of our model can be further improved by reducing the time steps to 2. In such case, our model still maintains superior performance. This demonstrates that our method has a better trade-off between the energy-efficiency and accuracy.
@@ Acknowledgements
This work was supported in part by the National Key R&D Program of China (Grant No. 2024YFB3311602); the National Natural Science Foundation of China (Grants 61971178, 62501224, 62272144, and U20A20183); the Ordos Science and Technology Major “Open Bidding” Project (Grant No. JBGS-2023-002); the State Grid Co., Ltd. Headquarters Technology Project (Grant No. 5500-202140127); the Major Project of Anhui Province (Grant No. 202423k09020001); the Anhui Provincial Natural Science Foundation for Distinguished Young Scholars (Grant No. 2408085J040); the Fundamental Research Funds for the Central Universities (Grants JZ2024HGTG0309, JZ2024AHST0337, JZ2024AKKG0507); and the Anhui Provincial Water Conservancy Science and Technology Project (Grant No. slky202501-06).
@@
CIFAR10. (krizhevsky2009learning) is a widely used dataset for image classification, consisting of 60,000 32x32 color images in 10 classes, with 6,000 images per class. The dataset is divided into 50,000 training images and 10,000 test images. In addition, we apply data augmentation techniques including data normalization, random horizontal flipping, random cropping, AutoAugment (cubuk2019autoaugment), and Cutout (devries2017improved), consistent with prior works (guo2024enof).
CIFAR100. (krizhevsky2009learning) is similar to CIFAR10 but contains 100 classes, each with 600 images. The dataset is also split into 50,000 training images and 10,000 test images. The data augmentation methods are consistent with those used for CIFAR10.
ImageNet. (deng2009imagenet) is a large-scale dataset for image classification, containing over 14 million images across 1,000 classes. The dataset is divided into training and validation sets, with 1.2 million training images and 50,000 validation images.
CIFAR10-DVS. (li2017cifar10) is a neuromorphic version of the CIFAR-100 dataset, containing 10 classes with 10,000 images. We follow the principles of prior works (wu2019direct) to split the dataset into 9,000 training images and 1,000 test images, resizing the images to 48x48 for model evaluation. In addition, we adopt the data augmentation strategy from (guo2022recdis), which includes random horizontal flipping and random cropping within 5 pixels.
@@
In our experiments, we use 8 NVIDIA 3090 GPUs for training on the ImageNet dataset, setting the batch size to 64 and the initial learning rate to 0.1. We employ the SGD optimizer with momentum set to 0.9 and weight decay of 1e-4, along with a Cosine Annealing learning rate scheduler. For the CIFAR10 and CIFAR100 datasets, we utilize 2 NVIDIA 3090 GPUs with a batch size of 128, keeping the other parameters consistent with those used for prior works (guo2024enof; xu2025bkdsnn). For the CIFAR100-DVS neuromorphic dataset, we use a single NVIDIA 3090 GPU and follow the EnoFSNN principles (guo2024enof) to adjust the first layer channel count of the ANN to 20, allowing it to process all time inputs at once. All models are implemented using PyTorch and PyTorch-Lightning frameworks.
@@
Maximum Entropy Principle is a probabilistic modeling method in information theory (jaynes1957information), which aims to select the probability distribution with maximum entropy under given constraints. The entropy is defined as:
@@
the first term $\sum_{k=1}^{K}q^{(k)}\log q^{(k)}$ is the negative value of the entropy $H(q)$, and maximizing $H(q)$ is equivalent to minimizing $\sum_{k=1}^{K}q^{(k)}\log q^{(k)}$. the second term $\lambda\left(\sum_{k=1}^{K}q^{(k)}-1\right)$ introduces the Lagrange multiplier $\lambda$ to handle the probability normalization constraint $\sum_{k=1}^{K}q^{(k)}=1$. According to the method of Lagrange multipliers, we can derive that the softmax is the unique solution to the maximum entropy distribution, but its specific form depends on the constraints (sun2024logit). When only the probability normalization constraint is present, the solution is a uniform distribution:
@@
Figure 10: Visualization of 2D loss landscape produced by different methods. Compared to without using KD or using prior KD methods (i.e., KDSNN (Xu2023kdsnn) and BKDSNN (xu2025bkdsnn)), student model trained with our CKDSNN has flatter loss landscape with fewer saddle points, leading to a smoother optimization path toward global minima.
@@
Applying SAMD in different stages. Teacher and student architectures, i.e., the ResNets and Transformer-based models, usually have four stages. We conduct experiments to explore the effects of applying our SAMD strategy in different stages. Tab. 7 reports the best performance when using SAMD at the final stage. This aligns with the convention of prior methods (xu2025bkdsnn; kim2021visual) to extract the class activation map (CAM) and spike activation map (SAM) at the final stage. The results show that CAM and SAM from deeper layers contain more precise semantic information related to the target class, resulting in a more effective distillation.
@@
Loss Landscape Visualization. In Fig. 10, we visualize the loss landscape of different KD methods. As can be seen, our method CKDSNN has a much flatter loss landscape compared with both the model trained without (w/o) KD and prior KD methods such as KDSNN (Xu2023kdsnn) and BKDSNN (xu2025bkdsnn). This indicates that our model can avoid the sharp convergence and gradient oscillation issues typically encountered around local minima during the training process, thereby guaranteeing more effective model learning. This should be thanks to the proposed noise-smoothed logits distillation strategy, supporting the student SNN model explore a smoother decision space to converge to a flatter local minima.

### v2

@@ Introduction
Results on ImageNet-1K. The comparison results are presented in Tab. 2. The proposed CKDSNN continues to outperform previous SOTA methods using various teacher models, including ResNet-18, ResNet-34, and ResNet-50. Specifically, compared to the prior SOTA methods BKDSNN (Xu et al. 2024), our CKDSNN improves the top-1 accuracy by 1.32%, 1.81%, and 1.52% using the three types of network architectures, respectively. These results indicate that CKDSNN is effective on the large-scale dataset.
Results on CIFAR10-DVS. The neuromorphic dataset comparison results are shown in Tab. 4. Our CKDSNN also outperforms existing methods on the neuromorphic dataset. For example, under the same architecture and time step settings, CKDSNN achieves an accuracy improvement of 1.05% compared to the most competitive EnOFSNN (Guo et al. 2024).
@@
STBP (Wu et al. 2019) NIPS’21 ResNet19 4 67.80
SEW-R (Fang et al. 2021) NIPS’21 WideNet 16 74.40
KDSNN (Xu et al. 2023) CVPR’23 ResNet20 10 78.31
BKDSNN (Xu et al. 2024) ECCV’24 ResNet20 10 79.53
EnOFSNN (Guo et al. 2024) NIPS’24 ResNet20 10 80.50
@@
Importance of CAM-SAM distillation in SAMD. The CAM-SAM distillation is one of the core components of SAMD. As show in Fig. 4 (b), when using ANN-based strategies (e.g., e2KD (Parchami-Araghi et al. 2024) or CATKD (Guo et al. 2023c)) to distill SNNs, the performance is significantly lower than our proposed CAM-SAM distillation. Specifically, CAM-CAM methods use class Grad-CAM strategies to generate activation maps, but the gradient estimation error leads to performance degradation. Additionally, CATKD is only applicable to CNN-base architectures. In contrast, SAM is designed on the characteristics of SNNs, enabling compatibility with various architectures.
@@
Figure 7: The t-SNE (Van der Maaten and Hinton 2008) visualization of feature distributions. (a): teacher ANN’s float feature representation. (b)-(d): spike features of student SNNs. Each color denotes an image class.
@@
Spiking Features. We extract spiking features from the final layer of our CKDSNN model and visualize them using t-SNE (Van der Maaten and Hinton 2008). Image samples from the CIFAR-100 dataset are used. As shown in Fig. 7, the spike features generated by CKDSNN demonstrate improved separability and discriminability compared to other KD methods and are close to the teacher’s feature distribution. This superiority is largely contributed to the proposed activation map KD, which guides the spiking representations to focus on salient regions, resulting in better feature discrimination.
@@
SEW-R (Fang et al. 2021) 18.0 4.14 4.03 67.04 4
KDSNN (Xu et al. 2023) 16.0 4.13 4.01 67.18 4
BKDSNN (Xu et al. 2024) 15.0 4.02 3.98 71.24 4
@@
Energy Efficiency. We finally evaluate the energy efficiency of our method by calculating the Fire rate, SOPS, and Power consumption following prior methods (Zhou et al. 2023; Xu et al. 2024; Fang et al. 2023). As shown in Tab. 6, compared with prior KD methods, our CKDSNN not only achieves state-of-the-art performance but offers higher energy efficiency. For example, our method with the same 4 time steps outperforms BKDSNN (Xu et al. 2024) by 1.81% but using slightly less fire rate and Power. The energy efficiency of our model can be further improved by reducing the time steps to 2. In such case, our model still maintains superior performance. This demonstrates that our method has a better trade-off between the energy-efficiency and accuracy.
@@ Acknowledgements
This work was supported in part by the National Key R&D Program of China (Grant No. 2024YFB3311602); the National Natural Science Foundation of China (Grants 61971178, 62501224, 62272144, and U20A20183); the Ordos Science and Technology Major “Open Bidding” Project (Grant No. JBGS-2023-002); the State Grid Co., Ltd. Headquarters Technology Project (Grant No. 5500-202140127); the Major Project of Anhui Province (Grant No. 202423k09020001); the Anhui Provincial Natural Science Foundation for Distinguished Young Scholars (Grant No. 2408085J040); the Fundamental Research Funds for the Central Universities (Grants JZ2024HGTG0309, JZ2024AHST0337, JZ2024AKKG0507); the Fundamental Research Funds for the Central Universities (JZ2025HGTA0160,JZ2025HGQA0139) and the Anhui Provincial Water Conservancy Science and Technology Project (Grant No. slky202501-06).
@@
CIFAR10. (Krizhevsky and Hinton 2009) is a widely used dataset for image classification, consisting of 60,000 32x32 color images in 10 classes, with 6,000 images per class. The dataset is divided into 50,000 training images and 10,000 test images. In addition, we apply data augmentation techniques including data normalization, random horizontal flipping, random cropping, AutoAugment (Cubuk et al. 2019), and Cutout (DeVries and Taylor 2017), consistent with prior works (Guo et al. 2024).
CIFAR100. (Krizhevsky and Hinton 2009) is similar to CIFAR10 but contains 100 classes, each with 600 images. The dataset is also split into 50,000 training images and 10,000 test images. The data augmentation methods are consistent with those used for CIFAR10.
ImageNet. (Deng et al. 2009) is a large-scale dataset for image classification, containing over 14 million images across 1,000 classes. The dataset is divided into training and validation sets, with 1.2 million training images and 50,000 validation images.
CIFAR10-DVS. (Li et al. 2017) is a neuromorphic version of the CIFAR-100 dataset, containing 10 classes with 10,000 images. We follow the principles of prior works (Wu et al. 2019) to split the dataset into 9,000 training images and 1,000 test images, resizing the images to 48x48 for model evaluation. In addition, we adopt the data augmentation strategy from (Guo et al. 2022), which includes random horizontal flipping and random cropping within 5 pixels.
@@
In our experiments, we use 8 NVIDIA 3090 GPUs for training on the ImageNet dataset, setting the batch size to 64 and the initial learning rate to 0.1. We employ the SGD optimizer with momentum set to 0.9 and weight decay of 1e-4, along with a Cosine Annealing learning rate scheduler. For the CIFAR10 and CIFAR100 datasets, we utilize 2 NVIDIA 3090 GPUs with a batch size of 128, keeping the other parameters consistent with those used for prior works (Guo et al. 2024; Xu et al. 2024). For the CIFAR100-DVS neuromorphic dataset, we use a single NVIDIA 3090 GPU and follow the EnoFSNN principles (Guo et al. 2024) to adjust the first layer channel count of the ANN to 20, allowing it to process all time inputs at once. All models are implemented using PyTorch and PyTorch-Lightning frameworks.
@@
Maximum Entropy Principle is a probabilistic modeling method in information theory (Jaynes 1957), which aims to select the probability distribution with maximum entropy under given constraints. The entropy is defined as:
@@
the first term $\sum_{k=1}^{K}q^{(k)}\log q^{(k)}$ is the negative value of the entropy $H(q)$, and maximizing $H(q)$ is equivalent to minimizing $\sum_{k=1}^{K}q^{(k)}\log q^{(k)}$. the second term $\lambda\left(\sum_{k=1}^{K}q^{(k)}-1\right)$ introduces the Lagrange multiplier $\lambda$ to handle the probability normalization constraint $\sum_{k=1}^{K}q^{(k)}=1$. According to the method of Lagrange multipliers, we can derive that the softmax is the unique solution to the maximum entropy distribution, but its specific form depends on the constraints (Sun et al. 2024). When only the probability normalization constraint is present, the solution is a uniform distribution:
@@
Figure 10: Visualization of 2D loss landscape produced by different methods. Compared to without using KD or using prior KD methods (i.e., KDSNN (Xu et al. 2023) and BKDSNN (Xu et al. 2024)), student model trained with our CKDSNN has flatter loss landscape with fewer saddle points, leading to a smoother optimization path toward global minima.
@@
Applying SAMD in different stages. Teacher and student architectures, i.e., the ResNets and Transformer-based models, usually have four stages. We conduct experiments to explore the effects of applying our SAMD strategy in different stages. Tab. 7 reports the best performance when using SAMD at the final stage. This aligns with the convention of prior methods (Xu et al. 2024; Kim and Panda 2021) to extract the class activation map (CAM) and spike activation map (SAM) at the final stage. The results show that CAM and SAM from deeper layers contain more precise semantic information related to the target class, resulting in a more effective distillation.
@@
Loss Landscape Visualization. In Fig. 10, we visualize the loss landscape of different KD methods. As can be seen, our method CKDSNN has a much flatter loss landscape compared with both the model trained without (w/o) KD and prior KD methods such as KDSNN (Xu et al. 2023) and BKDSNN (Xu et al. 2024). This indicates that our model can avoid the sharp convergence and gradient oscillation issues typically encountered around local minima during the training process, thereby guaranteeing more effective model learning. This should be thanks to the proposed noise-smoothed logits distillation strategy, supporting the student SNN model explore a smoother decision space to converge to a flatter local minima.